OpenStudy (anonymous):

Find the Taylor polynomials of orders 0, 1, 2, and 3 generated by f at x = a. f(x) = 1/x, a = 2

OpenStudy (anonymous):

I don't really understand the steps for solving these problems..

OpenStudy (experimentx):

f(2) = 1/2 f'(2) = -1/2^2 = -1/4 f'(3) = -1*-2/2^3 = 2/8 f'(4) = -1*-2*-3/16 = -6/16 so the taylor polynomial is f(x) = 1/2 - 1/4(x-2) + 2/8 (x-4)^2 - 6/16(x-3)^3 + O(4) something wrong http://www.wolframalpha.com/input/?i=taylor+expansion+1%2Fx+at+x%3D2

OpenStudy (amistre64):

why did you change a from 2 to 3 and 4?

OpenStudy (experimentx):

Oh ... forgot to divide by factorial :(

OpenStudy (anonymous):

wait could you explain how you got that :/ the process is a bit more important to me at the moment.

OpenStudy (amistre64):

the coeffs are generated by the first 3 derivatives of 1/x in this case evaluated at x=a

OpenStudy (amistre64):

then plug those into the general setup of the taylor

OpenStudy (experimentx):

\[ f(x) = \frac 1/2 - \frac 1/4(x-2 \frac 1{1!} + \frac 2/8 (x-4)^2 \frac 1{2!} - \frac 6/16(x-3)^3 \frac{1}{3!} + O(4) \] = Same answer was wolframalpha http://mathworld.wolfram.com/TaylorSeries.html Gotta go!!

OpenStudy (blockcolder):

You could just use the series for \(\frac{1}{1-x}\) and replace x with 1-x.

OpenStudy (amistre64):

thats the one i was thinking of ... maybe lol

OpenStudy (blockcolder):

Oh wait, it's supposed to be centered at a=2... nvm.

OpenStudy (anonymous):

I'm so confused.. our teacher decided to let a student from a BC calc class come teach our AB class, so the kid taught this to us in about 40 minutes. We're all really lost and I don't have much knowledge of taylor series or the way it works

OpenStudy (amistre64):

it works by constructing a polynomial to bend in the same way that the other function does

OpenStudy (amistre64):

we see how things bend by looking at their derivatives

OpenStudy (amistre64):

whats the derivative of 1/x?

OpenStudy (anonymous):

-1/x^2

OpenStudy (amistre64):

good, and the second derivative is?

OpenStudy (anonymous):

2/x^3

OpenStudy (amistre64):

great, and the 3rd derivative, since we want 1,2, and 3

OpenStudy (anonymous):

-6/(x^4)

OpenStudy (amistre64):

excellent, that gives us 4 numbers that we need to use; evaluate them at x=a=2

OpenStudy (anonymous):

so.. we're plugging 2 into all the derivatives?

OpenStudy (amistre64):

yes, in order to find out how the thing bends at x=2, we need to plug in x=2

OpenStudy (anonymous):

okay so we have 1/2, -1/4, 1/4, and -3/8

OpenStudy (amistre64):

i agree with this figures; now the next thing to do is to either try to remember the general setup of a taylor poly, or if you got time, we can see how one takes shape.

OpenStudy (amistre64):

the general set up is: \[T_{poly}=\frac{f^{(0)}(a)}{0!}(x-a)^0+\frac{f^{(1)}(a)}{1!}(x-a)^1+\frac{f^{(2)}(a)}{2!}(x-a)^2+\frac{f^{(3)}(a)}{3!}(x-a)^3...\]

OpenStudy (amistre64):

where those numerators are the numbers we determined just a few seconds ago

OpenStudy (anonymous):

what happens to the x? (in the (x-a))

OpenStudy (amistre64):

nothing, thats apart of the polynomial itself

OpenStudy (amistre64):

its like asking what happens to x in: y=2x

OpenStudy (anonymous):

ah so we don't have to plug anything in to it right

OpenStudy (amistre64):

its the independant variable that receives input in order to graph the thing

OpenStudy (amistre64):

well, not "as is" :)

OpenStudy (amistre64):

if you wanted to plot the poly, youd have to plug in values for x

OpenStudy (anonymous):

hmm. so how does this translate to the answers in my book? it says: P0(x) = 1/2 P1(x) = (1/2) - ((x-2)/4) P2(x) = (1/2) - ((x-2)/4) + (((x-2)^2)/8) P3(x) = (1/2) - ((x-2)/4) + (((x-2)^2)/8) - (((x-2)^3)/16)

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

nevermind that question

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

(I can't do simple math)

OpenStudy (anonymous):

Thank you so much for helping me! I really appreciated it!!

OpenStudy (amistre64):

youre welcome, and good luck :)